# The $\mathbb{S}_{n}$-equivariant Chow polynomial of the Braid matroid

### v2

@@ 2.3. The stratification by level trees
for the locally closed stratum where the level tree underlying the stable curve is equal to $(\mathbf{T},\ell)$.
@@
$\mathcal{B}_{(\mathbf{T},\ell)}\cong\prod_{j=0}^{k}\left(\prod_{v\in V_{j}(\mathbf{T})}\mathcal{M}_{0,\deg(v)}\times\mathbb{C}^{\star}\right)/\mathbb{C}^{\star},$
where the $\mathbb{C}^{\star}$-action on $\prod_{v\in V_{j}(\mathbf{T})}\mathcal{M}_{0,\operatorname{val}(v)}\times\mathbb{C}^{\star}$ is diagonal.
Proof.
The factors of $\mathcal{M}_{0,\mathrm{deg}(v)}$ correspond to the moduli space of choices for irreducible pointed curves at each vertex of the level tree. The additional factor of $\mathbb{C}^{\star}$ at a vertex $v$ corresponds to the choice of nonzero section of $\Omega_{C_{v}}(2)\cong\mathcal{O}_{\mathbb{P}^{1}}$, which determines a $\mathbb{C}^{\star}$-bundle over $\mathcal{M}_{0,\mathrm{deg}(v)}$. Since $\mathrm{Pic}(\mathcal{M}_{0,n})=0$ for all $n\geq 3$, this $\mathbb{C}^{\star}$-bundle must be trivial. The level-by-level quotient by a diagonal $\mathbb{C}^{\star}$-action corresponds to rescaling the differentials at a fixed level. ∎
@@ 3. Symmetric functions and $\mathbb{S}$-spaces
In this section we set up a motivic identity (Theorem 3.2), and explain how Theorem A follows from it. The proof of the identity will be given in §4. The identity is stated in the language of Serre characteristics of $\mathbb{S}$-spaces as developed by Getzler–Pandharipande [GP06], which we now briefly recall.
@@ 3.1. $\mathbb{S}$-spaces and associated operations
$(\mathcal{X}\boxtimes\mathcal{Y})(n)=\coprod_{k=0}^{n}\operatorname{Ind}_{\mathbb{S}_{k}\times\mathbb{S}_{n-k}}^{\mathbb{S}_{n}}\mathcal{X}(k)\times\mathcal{Y}(n-k).$
@@
where $\mathbb{S}_{k}$ acts diagonally.
@@ 3.2. From $\mathbb{S}$-spaces to symmetric functions
Let
$\Lambda:=\mathbb{Q}[\![p_{1},p_{2},\ldots]\!]$
denote the degree-completed ring of symmetric functions over $\mathbb{Q}$. Given a partition $\mu=(1^{\mu_{1}},2^{\mu_{2}},\ldots)$, we define the monomial
@@
$\operatorname{ch}_{n}(V):=\frac{1}{n!}\sum_{\tau\in\mathbb{S}_{n}}\mathrm{Tr}(\tau|V)\cdot p_{\lambda(\tau)}.$
@@
is a ring homomorphism which takes the composition of $\mathbb{S}$-spaces (3.1) into plethysm of symmetric functions [GP06, §5]. Plethysm is the operation
$\Lambda[\![t]\!]\times t\cdot\Lambda[\![t]\!]\to\Lambda[\![t]\!]$
@@
    for any $f\in t\cdot\Lambda[\![t]\!]$, the map $\Lambda[\![t]\!]\to\Lambda[\![t]\!]$ by $g\mapsto g\circ f$ is a $\mathbb{Q}$-algebra homomorphism;
@@
    if $f=f(t,p_{1},p_{2},\ldots)\in t\cdot\Lambda[\![t]\!]$, we have $p_{n}\circ f=f(t^{n},p_{n},p_{2n},\ldots)$.
@@ 3.3. A relation in $K_{0}(\mathsf{Var};\mathbb{S})$
We now define some important $\mathbb{S}$-spaces used in our proof of Theorem A. Let $D\mathcal{M}$ denote the $\mathbb{S}$-space
$D\mathcal{M}(n)=\begin{cases}\varnothing&\text{if }n\leq 1\\ \operatorname{Res}^{\mathbb{S}_{n+1}}_{\mathbb{S}_{n}}\mathcal{M}_{0,n+1}&\text{if }n\geq 2.\end{cases}$
Concretely, we think of $D\mathcal{M}(n)$ as the moduli space of smooth connected genus zero curves, together with $n+1$ ordered marked points $q_{0},\ldots,q_{n}$; the $\mathbb{S}_{n}$-action is by permutations of the last $n$ points. We also define an $\mathbb{S}$-space $\mathcal{B}$ by
@@
We also set the notation $\sigma_{n}$ for the class of $\operatorname{Spec}\mathbb{C}$ with trivial $\mathbb{S}_{n}$-action, considered as an $\mathbb{S}$-space supported in degree $n$. Finally, we will also consider $\mathbb{C}^{\star}$ as a $\mathbb{S}$-space supported in degree 0. Theorem A will be deduced from the following identity in $K_{0}(\mathsf{Var};\mathbb{S})$.
@@
$\mathcal{B}]\circ(\sigma_{1}+[\mathbb{C}^{\star}][D\mathcal{M}])=[\mathcal{B}]+[\mathbb{C}^{\star}]\cdot([\mathcal{B}]-[D\mathcal{M}])$
@@
By [Dev24, Proposition 13], we have $\operatorname{CH}^{k}(\mathcal{B}(n))\cong H^{2k}(\mathcal{B}(n);\mathbb{Q})$ for all $k$, and the space $\mathcal{B}(n)$ does not have any odd cohomology. Since $\mathcal{B}(n)$ is a smooth projective variety, $H^{2k}(\mathcal{B}(n);\mathbb{Q})$ is pure of weight $2k$. Therefore, we can write
@@
Moreover, we have $\mathsf{e}(\mathbb{C}^{\star})=t^{2}-1$, and
$\mathsf{e}(D\mathcal{M})=\mathsf{M}|_{t\mapsto t^{2}},$
@@
This holds because $H^{i}_{c}(\mathcal{M}_{0,n+1})$ is pure of weight 2(in+2)2(i-n+2), since it is a (n2)(n-2)-dimensional complex hyperplane arrangement complement. Making the relevant substitutions into Theorem 3.2, we find that
@@
To prove Theorem 3.2, we use the stratification of $\mathcal{B}_{n}$ by level trees. Recall that $\Gamma_{0,n}^{\mathrm{lev},\star}$ denotes the set of isomorphism classes of $(n+1)$-pointed stable level trees, and $\mathbb{S}_{n}$ naturally acts on this set by permuting the markings $\{1,\ldots,n\}$. Write
$\widetilde{\Gamma}_{0,n}^{\mathrm{lev},\star}:=\Gamma_{0,n}^{\mathrm{lev},\star}/\mathbb{S}_{n},$
and use the notation $\mathbf{T},\ell]$ for the $\mathbb{S}_{n}$-orbit of a level tree $(\mathbf{T},\ell)$. We define $\mathcal{B}_{[\mathbf{T},\ell]}$ to be an $\mathbb{S}$-space supported in degree $n$, where
@@ 4.1. Pruning extremal vertices
Given a level tree $(\mathbf{T},\ell)$, we write $V^{\mathrm{ext}}(\mathbf{T})$ for the set of vertices of $\mathbf{T}$ of the highest level, and set
@@
We define
@@
by deleting the $r(\mathbf{T},\ell)$ vertices of maximal depth from $\mathbf{T}$ and replacing them with markings; the ordering of these markings is not important, but to make a canonical choice we can use the lexicographic order on the subsets of $n$. The level structure $\ell^{\mathrm{pr}}$ is the restriction of the level structure $\ell$. See Figure 1 for an example. We call $(\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}})$ the pruning of $(\mathbf{T},\ell)$.
Figure 1. A level tree with 33 levels on the left, and its pruning on the right.
Now let
$U([\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}])\subseteq\coprod_{n\geq 2}\widetilde{\Gamma}_{0,n}^{\mathrm{lev,\star}}$
be the collection of all trees whose pruning is in the same $\mathbb{S}_{n}$-orbit as $\mathbf{T}^{\mathrm{pr}}$.
@@
We have
$\mathcal{B}_{[\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]}]\circ(\sigma_{1}+[\mathbb{C}^{\star}][D\mathcal{M}])=[\mathcal{B}_{[\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]}]+[\mathbb{C}^{\star}]\cdot\sum_{[\mathbf{T},\ell]\in U([\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}])}[\mathcal{B}_{[\mathbf{T},\ell]}].$
@@
Let $\mathcal{X}$ denote the $\mathbb{S}$-space defined by setting $\mathcal{X}(0)=\varnothing$ and
$\mathcal{X}(n)=\begin{cases}\sigma_{1}&\text{ if }n=1\\ \mathbb{C}^{\star}\times\operatorname{Res}^{\mathbb{S}_{n+1}}_{\mathbb{S}_{n}}\mathcal{M}_{0,n+1}&\text{ if }n\geq 2.\end{cases}$
Then a careful analysis of Lemma 2.9 yields
(4.2) $\coprod_{[\mathbf{T},\ell]\in U([\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}])}\mathcal{B}_{[\mathbf{T},\ell]}(n)\cong\left[\left(\mathcal{B}_{[\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]}\left(r(\mathbf{T},\ell)\right)\times\coprod_{\begin{subarray}{c}{i_{1}+\cdots+i_{r(\mathbf{T},\ell)}=n}\\ {\text{at least one }i_{j}>1}\end{subarray}}\operatorname{Ind}_{\prod_{j}\mathbb{S}_{i_{j}}}^{\mathbb{S}_{n}}\prod_{j=1}^{r(\mathbf{T},\ell)}\mathcal{X}(i_{j})\right)/\mathbb{S}_{r(\mathbf{T},\ell)}\right]/\mathbb{C}^{\star},$
where the $\mathbb{C}^{\star}$-action is free. The combinatorial insight is that a level tree in $U([\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}])$ is obtained by gluing a vertex to each marking of $\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]$: to make this precise, we choose an ordering of the $r(\mathbf{T},\ell)$ markings, and $r(\mathbf{T},\ell)$ ordered vertices $v_{i_{1}},\dots,v_{i_{r(\mathbf{T})}}$, each with $i_{j}$ markings, and then glue these vertices in the prescribed order to the markings. On the curve side, the terms where some $i_{j}=1$ can be visualized as vertices of valence $2$, and they correspond to cases where we decline to attach an irreducible component to the marking. The quotient by $\mathbb{S}_{r(\mathbf{T},\ell)}$ corresponds to compatibly reordering both the $r(\mathbf{T},\ell)$ vertices and the $r(\mathbf{T},\ell)$ markings. The quotient by the $\mathbb{C}^{\star}$-action corresponds to rescaling the differentials at the highest level. The condition that at least one $i_{j}>1$ reflects the condition that that the resulting tree is not equal to $\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]$. Since the $\mathbb{C}^{\star}$-action is free, (4.2) yields the following equality in the Grothendieck group of $\mathbb{S}_{n}$-varieties:
[…]
Now using the definition of composition of $\mathbb{S}$-spaces in (3.1), we see that
$\left[\left(\mathcal{B}_{[\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]}\left(r(\mathbf{T},\ell)\right)\times\coprod_{\begin{subarray}{c}{i_{1}+\cdots+i_{r(\mathbf{T},\ell)}=n}\\ {\text{at least one }i_{j}>1}\end{subarray}}\operatorname{Ind}_{\prod_{j}\mathbb{S}_{i_{j}}}^{\mathbb{S}_{n}}\prod_{j=1}^{r(\mathbf{T},\ell)}\mathcal{X}(i_{j})\right)/\mathbb{S}_{r(\mathbf{T},\ell)}\right]=[\mathcal{B}_{[\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]}]\circ[\mathcal{X}]-[\mathcal{B}_{[\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]}],$
so
$\mathbb{C}^{\star}]\cdot\sum_{[\mathbf{T},\ell]\in U([\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}])}[\mathcal{B}_{[\mathbf{T},\ell]}]=[\mathcal{B}_{[\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]}]\circ[\mathcal{X}]-[\mathcal{B}_{[\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]}].$
After substituting $\mathcal{X}]=\sigma_{1}+[\mathbb{C}^{\star}][D\mathcal{M}]$ to the right-hand side of the above equality and adding $\mathcal{B}_{[\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]}]$ to both sides, we obtain the statement of the lemma. ∎
@@
Let $\mathcal{B}^{(k)}\subset\mathcal{B}$ denote the $\mathbb{S}$-subspace parameterizing stable curves whose level tree has length $k$. Then by summing the equality of Lemma 4.1 over all $\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}}]$ with length equal to $k$, we obtain
$\mathcal{B}^{(k)}]\circ(\sigma_{1}+[\mathbb{C}^{\star}][D\mathcal{M}])=[\mathcal{B}^{(k)}]+[\mathbb{C}^{\star}][\mathcal{B}^{(k+1)}]$
for $k\geq 1$. Summing over all $k\geq 1$, we see that
$\mathcal{B}]\circ(\sigma_{1}+[\mathbb{C}^{\star}][D\mathcal{M}])=[\mathcal{B}]+[\mathbb{C}^{\star}]([\mathcal{B}]-[\mathcal{B}^{(1)}]).$
Theorem 3.2 now follows from the fact that $\mathcal{B}^{(1)}]=[D\mathcal{M}]$. ∎

### v3

@@ 2.4. The stratification by level trees
for the locally closed stratum where the level tree underlying the stable curve is equal to $(\mathbf{T},\ell)$. The following lemma is immediate from the definition of a multiscale differential.
@@
$\mathcal{B}_{(\mathbf{T},\ell)}\cong\prod_{j=0}^{k}\left(\prod_{v\in V_{j}(\mathbf{T})}\Omega^{\star}\mathcal{M}_{0,\deg(v)}\right)/\mathbb{C}^{\star},$
where the $\mathbb{C}^{\star}$-action on $\prod_{v\in V_{j}(\mathbf{T})}\Omega^{\star}\mathcal{M}_{0,\deg(v)}$ is diagonal and free.
@@ 3. Symmetric functions and $\mathbb{S}$-spaces
In this section we set up a motivic identity (Theorem 3.5), and explain how Theorem A follows from it. The proof of the identity will be given in §4. The identity is stated in the language of Serre characteristics of $\mathbb{S}$-spaces as developed by Getzler–Pandharipande [GP06], which we now recall.
@@ 3.1. $\mathbb{S}$-spaces and associated operations
(3.1) $(\mathcal{X}\boxtimes\mathcal{Y})(n)=\coprod_{k=0}^{n}\operatorname{Ind}_{\mathbb{S}_{k}\times\mathbb{S}_{n-k}}^{\mathbb{S}_{n}}\mathcal{X}(k)\times\mathcal{Y}(n-k).$
@@
where $\mathbb{S}_{k}$ acts diagonally. Note that the quotient by $\mathbb{S}_{k}$ exists as a variety since $\mathbb{S}_{k}$ is a finite group.
The following basic combinatorial lemma about compositions of $\mathbb{S}$-spaces is essentially immediate from the definition, but it will be a useful identity for a later proof.
Lemma 3.3.
Suppose $\mathcal{X}$ and $\mathcal{Y}$ are $\mathbb{S}$-spaces such that $\mathcal{Y}(n)=\varnothing$ for $n\leq 1$. Let
$\mathcal{Z}(n):=\begin{cases}\operatorname{Spec}\mathbb{C}&\text{if }n=1\\ \mathcal{Y}(n)&\text{if }n\neq 1.\end{cases}$
Then in $K_{0}(\mathsf{Var};\mathbb{S})$, we have
$\mathcal{X}\circ\mathcal{Z}]=[\mathcal{W}]$
where in $K_{0}(\mathsf{Var};\mathbb{S}_{n})$, we have
[…]
Proof.
From the definition of composition (3.2) we have the identity
$\mathcal{W}(n)]=\sum_{j=0}^{n}[(\mathcal{X}(j)\times\mathcal{Z}^{\boxtimes j}(n))/\mathbb{S}_{j}]$
in the Grothendieck group of $\mathbb{S}_{n}$-varieties. From the definition of $\mathcal{Z}$, we find that summand corresponding to $j=n$ is exactly $\mathcal{X}(n)]$. Now the formula follows by applying the definition of the box product (3.1) to each remaining term in the sum: if we have $i_{1}+\cdots+i_{j}=n$ for positive integers $i_{\ell}$, and $j<n$, then it is automatic that at least one $i_{\ell}>1$. ∎
@@ 3.2. From $\mathbb{S}$-spaces to symmetric functions
Define the graded ring
$\Lambda:=\mathbb{Q}[\![p_{1},p_{2},\ldots]\!],$
where $p_{i}$ has degree $i$. The ring $\Lambda$ is the completion, with respect to degree, of the usual ring of symmetric functions over $\mathbb{Q}$. Elements of $\Lambda$ are expressions of the form
$f=\sum_{d\geq 0}f_{d}$
where each $f_{d}=f_{d}(p_{1},p_{2},\ldots)$ is a polynomial which is homogeneous of degree $d$. Given a partition $\mu=(1^{\mu_{1}},2^{\mu_{2}},\ldots)$, we define the monomial
@@
(3.4) $\operatorname{ch}_{n}(V):=\frac{1}{n!}\sum_{\tau\in\mathbb{S}_{n}}\mathrm{Tr}(\tau|V)\cdot p_{\lambda(\tau)}.$
@@
is a ring homomorphism which takes the composition of $\mathbb{S}$-spaces (3.2) into plethysm of symmetric functions [GP06, §5], defined as follows. First, set
$F_{1}\Lambda[\![t]\!]\subset\Lambda[\![t]\!]$
to be the vector subspace of elements with no constant term; i.e. the terms which do not have bidegree $(0,0)$. Plethysm is the operation
$\Lambda[\![t]\!]\times F_{1}\Lambda[\![t]\!]\to\Lambda[\![t]\!]$
@@
    for any $f\in F_{1}\Lambda[\![t]\!]$, the map $\Lambda[\![t]\!]\to\Lambda[\![t]\!]$ by $g\mapsto g\circ f$ is a $\mathbb{Q}$-algebra homomorphism;
@@
    for any $n$ and $f\in F_{1}\Lambda[\![t]\!]$, we have $p_{n}\circ f=\psi_{n}(f)$, where
    $\psi_{n}:\Lambda[\![t]\!]\to\Lambda[\![t]\!]$
    is the algebra homomorphism defined by $t\mapsto t^{n}$ and $p_{k}\mapsto p_{kn}$ for all $k>0$.
@@ 3.3. A relation in $K_{0}(\mathsf{Var};\mathbb{S})$
We now define some important $\mathbb{S}$-spaces used in our proof of Theorem A. Let $\Omega^{\star}\mathcal{M}$ denote the $\mathbb{S}$-space defined by
$\Omega^{\star}\mathcal{M}(n)=\begin{cases}\varnothing&\text{if }n\leq 1\\ \Omega^{\star}\mathcal{M}_{0,n+1}&\text{if }n\geq 2\end{cases}$
(recall that $\Omega^{\star}\mathcal{M}_{0,n+1}$ was defined in Definition 2.9). We also define an $\mathbb{S}$-space $\mathcal{B}$ by
@@
We also set the notation $\sigma_{n}$ for the class of $\operatorname{Spec}\mathbb{C}$ with trivial $\mathbb{S}_{n}$-action, considered as an $\mathbb{S}$-space supported in degree $n$. Theorem A will be deduced from the following identity in $K_{0}(\mathsf{Var};\mathbb{S})$.
@@
$\mathcal{B}]\circ(\sigma_{1}+[\Omega^{\star}\mathcal{M}])=[\mathcal{B}]+[\mathbb{C}^{\star}][\mathcal{B}]-[\Omega^{\star}\mathcal{M}]$
@@
By [Dev26, Proposition 13], we have $\operatorname{CH}^{k}(\mathcal{B}(n))\cong H^{2k}(\mathcal{B}(n);\mathbb{Q})$ for all $k$, and the space $\mathcal{B}(n)$ does not have any odd cohomology. Since $\mathcal{B}(n)$ is a smooth projective variety, $H^{2k}(\mathcal{B}(n);\mathbb{Q})$ is pure of weight $2k$. Therefore, we can write
@@
Moreover, by Lemma 2.10, we have
$\mathsf{e}(\Omega^{\star}\mathcal{M})=\mathsf{e}(\mathbb{C}^{\star})\cdot\mathsf{M}|_{t\mapsto t^{2}},$
@@
This holds because $H^{i}_{c}(\mathcal{M}_{0,n+1})$ is pure of weight 2(in+2)2(i-n+2), since it is a (n2)(n-2)-dimensional complex hyperplane arrangement complement. Using $\mathsf{e}(\mathbb{C}^{\star})=t^{2}-1$ and making the relevant substitutions into Theorem 3.5, we find that
@@
To prove Theorem 3.5, we use the stratification of $\mathcal{B}_{n}$ by level trees. The basic idea is that the composition $\mathcal{B}]\circ(\sigma_{1}+[\Omega^{\star}\mathcal{M}])$ on the left-hand side of Theorem 3.5 reflects the combinatorial operation of adding either one or zero levels to a level tree, as we will make precise in this section.
Recall that $\Gamma_{0,n}^{\mathrm{lev},\star}$ denotes the set of isomorphism classes of $(n+1)$-pointed stable level trees, and $\mathbb{S}_{n}$ naturally acts on this set by permuting the markings $\{1,\ldots,n\}$.
Definition 4.1.
We use the notation $\mathbf{T},\ell]$ for the $\mathbb{S}_{n}$-orbit of a level tree $(\mathbf{T},\ell)$. We define $\mathcal{B}_{[\mathbf{T},\ell]}$ to be an $\mathbb{S}$-space supported in degree $n$, where
@@
Now we define another $\mathbb{S}$-space $\Omega^{\star}\mathcal{M}_{aug}$.
Definition 4.2.
Define $\Omega^{\star}\mathcal{M}_{aug}$ by
$\Omega^{\star}\mathcal{M}_{aug}(n)=\begin{cases}\operatorname{Spec}\mathbb{C}&\text{if }n=1\\ \Omega^{\star}\mathcal{M}(n)&\text{otherwise.}\end{cases}$
The reason for defining $\Omega^{\star}\mathcal{M}_{aug}$ is that in $K_{0}(\mathsf{Var};\mathbb{S})$, we have the identity
$\Omega^{\star}\mathcal{M}_{aug}]=\sigma_{1}+[\Omega^{\star}\mathcal{M}].$
4.1. Boxed powers of $\Omega^{\star}\mathcal{M}_{aug}$
Towards understanding the composition $\mathcal{B}\circ\Omega^{\star}\mathcal{M}_{aug}]$, we will first state how to think of each boxed power $\Omega^{\star}\mathcal{M}_{aug}^{\boxtimes r}$ as an auxiliary moduli space for each integer $r\geq 1$. Indeed, unwinding the definition of the box product (3.1), we find that for each $n\geq 1$, we have that $\Omega^{\star}\mathcal{M}_{aug}^{\boxtimes r}(n)$ is the moduli space of ordered tuples
$(C_{1},\ldots,C_{r},f)$
where
  1. (1)
    $C_{j}$ is either
    • a multiscale differential with distinguished marking $q_{j}$, as well as $i_{j}$ additional non-distinguished markings $p_{k_{1}},\ldots,p_{k_{i_{j}}}$, or
    • a single point $p_{j}$ which we may treat as a non-distinguished marking.
  2. (2)
    If we let $M$ denote the set $\{p\mid p\in C_{j}\,\text{a non-distinguished marking for some }j\}$, then
    $f:M\to\{1,\ldots,n\}$
    is a bijection.
The $\mathbb{S}_{n}$-action on $\Omega^{\star}\mathcal{M}_{aug}^{\boxtimes r}(n)$ is induced by permutations of $\{1,\ldots,n\}$. Note that there is a natural $\mathbb{S}_{r}$-action on $\Omega^{\star}\mathcal{M}^{\boxtimes r}_{aug}$ by re-ordering the multiscale differentials. This $\mathbb{S}_{r}$-action appears in the definition of the composition operation (3.2).
@@ 4.2. Pruning extremal vertices
We now turn towards a combinatorial interpretation of the composition $\mathcal{B}\circ\Omega^{\star}\mathcal{M}_{aug}]$. Given a level tree $(\mathbf{T},\ell)$, we write $V^{\mathrm{ext}}(\mathbf{T})$ for the set of vertices of $\mathbf{T}$ of the highest level, and set
@@
Definition 4.3.
For a level tree $(\mathbf{T},\ell)\in\Gamma_{0,n}^{\mathrm{lev},\star}$, we define the pruning of $(\mathbf{T},\ell)$ as a level tree
@@
by deleting the $r(\mathbf{T},\ell)$ vertices in $V^{\mathrm{ext}}(\mathbf{T})$ from $\mathbf{T}$ and replacing them with markings. The ordering of the markings on $(\mathbf{T}^{\mathrm{pr}},\ell^{\mathrm{pr}})$ is the unique one which is compatible with the lexicographic ordering of the subsets of $\{1,\ldots,n\}$. The level structure $\ell^{\mathrm{pr}}$ is the restriction of the level structure $\ell$.
Note that the pruning always has exactly one less level than the original level tree. See Figure 1 for an example of a pruning.
Figure 1. A level tree with 33 levels on the left, and its pruning on the right. Here, a marking on a vertex $v$ is visualized by a labelled half-edge emanating from $v$.
Now we define
$U([\mathbf{T},\ell])\subseteq\coprod_{k\geq 2}\Gamma_{0,k}^{\mathrm{lev,\star}}$
to be the set of all level trees whose pruning is in the same $\mathbb{S}_{n}$-orbit as $(\mathbf{T},\ell)$. The following lemma encodes the recursive structure obtained by pruning, at the level of $\mathbb{S}$-spaces.
@@
In the Grothendieck ring of $\mathbb{S}$-spaces, the identity
$\mathcal{B}_{[\mathbf{T},\ell]}]\circ(\sigma_{1}+[\Omega^{\star}\mathcal{M}])=[\mathcal{B}_{[\mathbf{T},\ell]}]+[\mathbb{C}^{\star}]\cdot\sum_{[\mathbf{T}^{\prime},\ell^{\prime}]\in U([\mathbf{T},\ell])}[\mathcal{B}_{[\mathbf{T}^{\prime},\ell^{\prime}]}]$
holds.
@@
The left-hand side is the class in $K_{0}(\mathsf{Var};\mathbb{S})$ of
$\mathcal{B}_{[\mathbf{T},\ell]}\circ\Omega^{\star}\mathcal{M}_{aug},$
where $\Omega^{\star}\mathcal{M}_{aug}$ is the $\mathbb{S}$-space defined in Definition 4.2. Then Lemma 3.3 implies that
(4.5) $\mathcal{B}_{[\mathbf{T},\ell]}\circ\Omega^{\star}\mathcal{M}_{aug}]=[\mathcal{B}_{[\mathbf{T},\ell]}]+\left[\mathcal{X}\right]$
where $\mathcal{X}$ is the $\mathbb{S}$-space such that
(4.6) $\mathcal{X}(n)=\left(\mathcal{B}_{[\mathbf{T},\ell]}\left(r(\mathbf{T},\ell)\right)\times\coprod_{\begin{subarray}{c}{i_{1}+\cdots+i_{r(\mathbf{T},\ell)}=n}\\ {\text{at least one }i_{j}>1}\end{subarray}}\operatorname{Ind}_{\prod_{j}\mathbb{S}_{i_{j}}}^{\mathbb{S}_{n}}\prod_{j=1}^{r(\mathbf{T},\ell)}\Omega^{\star}\mathcal{M}(i_{j})\right)/\mathbb{S}_{r(\mathbf{T},\ell)}$
The disjoint union appearing in (4.6) is a subspace of the moduli space $\Omega^{\star}\mathcal{M}_{aug}^{\boxtimes r(\mathbf{T},\ell)}(n)$ described in §4.1: it is exactly the subspace where at least one of the $C_{1},\ldots,C_{r(\mathbf{T},\ell)}$ is not a point (that is, we have at least one honest differential). Now generalizing Lemma 2.12, we find that
(4.7) $\mathcal{X}(n)/\mathbb{C}^{\star}=\coprod_{[\mathbf{T}^{\prime},\ell^{\prime}]\in U([\mathbf{T},\ell])}\mathcal{B}_{[\mathbf{T}^{\prime},\ell^{\prime}]}(n),$
where the disjoint union is finite, because only finitely many $\mathbf{T}^{\prime},\ell^{\prime}]\in U([\mathbf{T},\ell])$ can have $\mathcal{B}_{[\mathbf{T}^{\prime},\ell^{\prime}]}(n)\neq\varnothing$. Using the interpretation in §4.1, the right-hand side of (4.6) should be thought of as adding a single level to a multiscale differential in $\mathcal{B}_{[\mathbf{T},\ell]}$, and the (free) $\mathbb{C}^{\star}$-quotient on the left-hand side of (4.7) corresponds to rescaling the differentials at the new level. Considering all $n$ at once, we find
(4.8) $\mathcal{X}]=[\mathbb{C}^{\star}]\cdot\sum_{[\mathbf{T}^{\prime},\ell^{\prime}]\in U([\mathbf{T},\ell])}[\mathcal{B}_{[\mathbf{T}^{\prime},\ell^{\prime}]}$
in $K_{0}(\mathsf{Var};\mathbb{S})$. Substituting (4.8) into (4.5), we obtain the desired statement. ∎
@@
Let $\mathcal{B}^{(k)}\subset\mathcal{B}$ denote the $\mathbb{S}$-subspace parameterizing stable curves whose level tree has length $k$ (recall that the length of a level tree is the number of levels). Then by summing the equality of Lemma 4.4 over all $\mathbf{T},\ell]$ with length equal to $k$, we obtain
$\mathcal{B}^{(k)}]\circ(\sigma_{1}+[\Omega^{\star}\mathcal{M}])=[\mathcal{B}^{(k)}]+[\mathbb{C}^{\star}][\mathcal{B}^{(k+1)}]$
for $k\geq 1$. Now we sum over all $k\geq 1$ to find that
$\mathcal{B}]\circ(\sigma_{1}+[\Omega^{\star}\mathcal{M}])=[\mathcal{B}]+[\mathbb{C}^{\star}]([\mathcal{B}]-[\mathcal{B}^{(1)}]).$
Theorem 3.5 now follows from Lemma 2.10, which implies that $\mathbb{C}^{\star}][\mathcal{B}^{(1)}]=[\Omega^{\star}\mathcal{M}]$. ∎